Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-02-16

Observing: Stefan Taubenberger, Tamar Faran, Markus Kromer

Support Team: Nancy Elias-Rosa

TAT: Morgan Fraser, Lluís Galbany


Weather conditions


Start of night: some scattered clouds, humidity ~60%, wind ~1.5 m/s

Middle of night: thin cloud cover, humidity ~30%, wind ~ 6.5m/s

End of night: clear, humidity ~25%, wind ~ 8m/s


------------------------------------------------


Calibrations


Evening: PESSTO_SkyFF_eve_grz

Morning: -

------------------------------------------------



00:10 UT: data analysis

00:45 UT: Follow-up:  SN2016ado, gr16 2700s , seeing ~1.6" (the first set of internal flats was not good, repeated immediately)
01:50 UT: Follow-up:  OGLE15xl, gr13 3x2700s , seeing ~1.2" 

04:20 UT: Standard: LTT3864, gr11+gr13+gr16, seeing ~1" (more clouds have moved in, still clear around the horizon)
04:40 UT: classification: LSQ16lq - target could not be observed in the acquisition image

01:50 UT: classification:  LSQ16ol, gr13, 1500s, seeing ~1"

05:30 UT: Follow-up:  SN2016O, gr13, 2x1800s, seeing ~1"

06:35 UT: classification: PS16agp, gr13, 1500s, seeing ~1.4"

07:10 UT: classification: LSQ16on - target could not be observed in the acquisition image

07:15 UT: Follow-up:  SN2016adj, gr11 1000s, gr16 150s, seeing ~1.2"
08:00 UT: Follow-up: MASTERJ141023.42-431843.7, gr13 1800s, seeing ~1.1"


08:30 UT: End of the night - a problem with the temperature of the rotator, observations stopped




Comments